### v3

# arXiv:2503.21735v3[cs.SE] 01 Mar 2026

Abstract
Ensuring reliable data-driven decisions is crucial in domains where analytical accuracy directly impacts safety, compliance, or operational outcomes. Decision support in such domains relies on large tabular datasets, where manual analysis is slow, costly, and error-prone. While Large Language Models (LLMs) offer promising automation potential, they face challenges in analytical reasoning, structured data handling, and ambiguity resolution. This paper introduces GateLens, an LLM-based architecture for reliable analysis of complex tabular data. Its key innovation is the use of Relational Algebra (RA) as a formal intermediate representation between natural-language reasoning and executable code, addressing the reasoning-to-code gap that can arise in direct generation approaches. In our automotive instantiation, GateLens translates natural language queries into RA expressions and generates optimized Python code. Unlike traditional multi-agent or planning-based systems that can be slow, opaque, and costly to maintain, GateLens emphasizes speed, transparency, and reliability. We validate the architecture in automotive software release analytics, where experimental results show that GateLens outperforms the existing Chain-of-Thought (CoT) + Self-Consistency (SC) based system on real-world datasets, particularly in handling complex and ambiguous queries. Ablation studies confirm the essential role of the RA layer. Industrial deployment demonstrates over 80% reduction in analysis time while maintaining high accuracy across domain-specific tasks. GateLens operates effectively in zero-shot settings without requiring few-shot examples or agent orchestration. This work advances deployable LLM system design by identifying key architectural features—intermediate formal representations, execution efficiency, and low configuration overhead—crucial for domain-specific analytical applications where accuracy, traceability, and stakeholder trust are paramount.
keywords:
Large Language Models, Tabular Question Answering, Software Release Analytics, Automotive Software Testing, Test Result Interpretation, Interpretable Reasoning
journal: Journal of Systems and Software
[…]
[label1]organization=Department of Computer Science and Engineering, Chalmers University of Technology, city=Gothenburg, country=Sweden
[…]
[label2]organization=Volvo Group, city=Gothenburg, country=Sweden
1 Introduction
Reliable decision-making in data-intensive domains depends on the ability to accurately analyze large volumes of structured data. In sectors such as automotive manufacturing, healthcare, finance, and regulatory compliance, critical decisions hinge on interpreting tabular datasets that capture test results, operational metrics, or validation records. These datasets often pass through gating steps—critical checkpoints where predefined quality or compliance standards must be met. Failures at these gates can cascade through interconnected processes, delaying dependent workflows regardless of their individual quality. Analysts tasked with safeguarding decision quality must process vast quantities of data. While essential for ensuring accuracy and reliability, this process is time-consuming and prone to human error in data interpretation.
The software industry’s transition from manual to automated processes has entered a new era with the emergence of Large Language Models (LLMs) [10.3390/app14031046, chang2024survey]. Companies are increasingly integrating these AI agents into their workflows, seeking more cost-effective and optimized solutions for complex software engineering tasks [10.1109/ase56229.2023.00035]. However, direct application of LLMs to structured data analysis for decision support is hindered by limitations in interpretable reasoning and understanding of technical specifications [10.3390/fi16060180, austin2021program].
Figure 1: Comparison of conventional Chain-of-Thought (CoT) reasoning and its extension using Relational Algebra (RA) statements. (Left) CoT employs unstructured, fused reasoning where multiple analytical concepts are blended together in informal thoughts without clear separation. Operations cannot be independently mapped to executable code snippets, making reasoning steps opaque and difficult to debug in isolation. (Right) RA-based reasoning adopts a structured, compositional approach where each relational algebra operation (filtering, joining, projecting) is a discrete, formally grounded step. Each operation is independently interpretable, reusable, and can be directly mapped to executable code. The color-coding illustrates how RA maintains clear boundaries between distinct analytical steps, whereas CoT conflates multiple operations within single reasoning thoughts. This structural distinction enables transparent, formally grounded reasoning that can be systematically verified and debugged at each stage.
To address these challenges, we introduce GateLens, a reasoning-enhanced LLM agent [miehling2025agentic] for reliable tabular data analysis to support decision-making in domain-specific contexts. We validate the architecture in the automotive software release domain, where the need for precise, transparent, and efficient analysis is particularly acute.
A key challenge in LLM-based analytical agents is the potential mismatch between natural-language reasoning and the actual computation implemented in generated code—a reasoning-to-code gap that becomes more pronounced as analytical complexity increases. GateLens addresses this challenge by integrating structured relational analysis with domain-specific expertise through a reasoning layer built on Relational Algebra (RA) as an extension of Chain-of-Thought (CoT). This reasoning layer systematically breaks down complex validation tasks into discrete, formally grounded analytical steps. We adopt this approach to address essential limitations of vanilla CoT for our work: its reasoning steps are opaque and often cannot be mapped directly to executable code, its operations are not compositional and cannot be independently reused or debugged, and its reasoning is unstructured with operations blended together without clear intent or formal grounding. In contrast, our RA-based reasoning layer aims to ensure that each step is independently interpretable and reusable, reasoning is grounded in formal relational algebra, and intent is made explicit through well-defined operations. Figure 1 illustrates this distinction between our structured RA-based approach and conventional CoT reasoning.
GateLens simplifies three critical aspects of release validation:
1. Test Result Analysis: Analyzing test execution outcomes is foundational to release validation. This involves analyzing pass/fail patterns across comprehensive test suites, identifying recurring failures, and validating test coverage metrics. In automotive software, where a single release might involve a large number of test cases across multiple vehicle functions, this task becomes particularly demanding. Release engineers must not only identify failed tests but also understand their patterns, assess coverage adequacy, and evaluate test execution stability.
2. Impact Assessment: Impact Assessment is a systematic process for evaluating how software issues affect vehicle functionality and safety during release validation. It involves three phases: first, a critical failure analysis identifies the root cause and immediate effects of an issue, such as an ABS module causing a 200ms brake signal delay that exceeds the 100ms threshold. Second, a component-level impact evaluation traces how the issue propagates through interconnected systems, assessing both direct effects, like problematic emergency braking, and indirect effects, such as reduced stability control performance. Finally, an integration risk assessment quantifies the severity of these impacts against safety thresholds and functional requirements, categorizing issues like the ABS delay as system-wide risks with critical severity. This structured process enables engineers to understand system-wide effects, ensuring all safety and functionality requirements are met before release.
3. Release Candidate Analysis: The final quality gate involves evaluating Release Candidates (RCs) against predefined quality gates and criteria. This encompasses analyzing whether a particular RC meets all quality thresholds, identifying potential release blockers, and validating compliance with release requirements. In automotive software, where releases must meet stringent safety and quality standards, this analysis requires careful validation of each RC against established criteria, ensuring all prerequisites for a safe and reliable release are satisfied.
The traditional release validation process demands extensive manual effort. Release engineers meticulously analyze test results, assess impacts, verify RCs against quality gates, and report findings to stakeholders, such as release managers. As automotive software systems grow increasingly complex, these manual workflows become more challenging, time-consuming, and error-prone.
This work aims to streamline release validation by automating key analysis workflows, enabling engineers to focus on high-value analysis and discussion. By providing deeper analytical insights, the proposed approach reduces the time needed to deliver accurate validation results, empowering release managers to make informed decisions more efficiently. Our contributions can be summarized as follows:
  • We design an architecture optimized for time- and safety-critical environments, minimizing LLM invocations while preserving reasoning depth for reliable tabular analysis. GateLens operates in a zero-shot setting, avoiding the need for few-shot examples or multi-agent coordination, which improves generalizability, execution speed, reduces maintenance overhead, and enhances transparency.
  • We introduce a scalable and maintainable framework for automotive software release validation, developed in response to observed limitations in traditional planning-based multi-agent system. GateLens handles diverse user queries with higher robustness and clarity, supporting effective decision-making across a wider range of release engineering tasks.
  • We conduct a comprehensive empirical evaluation, including comparisons with a CoT+SC system, ablation of the RA module, and performance across multiple LLMs (GPT-4o and Llama 3.1 70B). These experiments demonstrate GateLens’s performance advantages in complex and ambiguous queries.
  • We report on real-world industrial deployment in a partner automotive company, where GateLens reduces analysis time by over 80% and demonstrates strong generalization across user roles, highlighting its practical value and deployment-readiness in safety-critical release validation workflows.
2 Background and Motivation
Software release decisions in the automotive industry involve multiple stakeholders and extensive data analysis. Modern vehicles integrate hundreds of software components, each requiring rigorous testing and validation. The process advances through distinct phases: component-level testing, integration testing, system-level validation, and vehicle validation testing. Component-level testing verifies individual software modules, integration testing ensures proper interaction between components, system-level validation examines the complete system behavior, and vehicle validation testing evaluates software performance under real vehicle conditions on closed tracks.
The development cycle grows in complexity with each integration phase. This increasing complexity presents challenges in managing large-scale test results, tracking interdependencies between components, correlating test failures across different subsystems, and maintaining historical context for recurring issues. The iterative nature of software testing and validation further expands this data ecosystem.
The wide range of stakeholders in the release process creates additional challenges in data interpretation and presentation. Project managers need high-level progress indicators, verification engineers require detailed technical insights, quality engineers focus on trend analysis and improvement metrics, and release engineers need specific release-readiness indicators. This variety of perspectives necessitates different views of the same underlying data, making the analysis process more complex.
This diversity is reflected not only in perspective but also in the level of granularity required from the underlying data. For example, senior management may ask high-level questions such as: “List all vehicles that have not yet received global approval,” seeking a consolidated overview without reference to schema details or subsystem-specific attributes. In contrast, release managers or verification engineers may pose highly specific queries such as: “What are the baseline, phase, RC, and EUF values for RM-320 in the latest test suite?” These questions rely on detailed knowledge of domain terminology and schema structure. Although both types of queries operate on the same data foundation, they demand substantially different views and levels of abstraction.
Release managers function as gatekeepers in the software deployment pipeline. They handle test result analysis, cross-system impact assessment, decision-making, stakeholder coordination, and safety compliance verification. The manual workflow introduces vulnerabilities: time-intensive processing, potential errors in interpretation, decision delays, and communication barriers between technical and business teams.
Within this process, statisticians provide an overall view of the data to project managers and quality engineers for future business decisions. The existing manual approach faces several limitations, particularly regarding time and resource constraints. These include labor-intensive data analysis, delayed response to critical issues, limited capacity for comprehensive analysis, and bottlenecks in the release pipeline. Communication challenges further complicate the process, with misalignment between technical analysts and statisticians, varying interpretations of project requirements, inconsistent reporting formats, and knowledge transfer gaps.
Internal Testing on Closed Track represents a crucial validation phase. Release managers must analyze extensive datasets to evaluate progression readiness. The manual query process for report generation can impact release timelines, business objectives, and subsystem integration schedules.
The deployment of intelligent assistants presents opportunities to address these challenges through automated data processing and analysis, standardized reporting frameworks, real-time insights generation, and stakeholder-specific view generation. However, current automation solutions, including basic LLM implementations, face limitations in understanding complex technical specifications, maintaining structured analysis steps, handling domain-specific requirements, and processing automotive validation data systematically.
These limitations highlight the need for enhanced solutions combining domain expertise with advanced analytical reasoning capabilities. Such solutions must facilitate efficient decision-making while maintaining high safety and quality standards in automotive software development [wang2024surveyl]. Effective intelligent assistants can transform the release decision process, enabling release managers to prioritize result interpretation and strategic decision-making over routine data analysis.
3 Approach and Methodology
The complexity of automotive software release validation demands a system that can bridge the gap between human-centric inquiries and precise technical analysis. GateLens addresses this challenge by utilizing LLM agents to transform natural language queries into actionable insights through systematic analysis. At its core, GateLens must fulfill three fundamental requirements:
1. Query Understanding: The system must accurately interpret diverse user queries, ranging from high-level management questions to detailed technical inquiries about specific test cases.
2. Query Transformation: The system needs to transform these interpretations into structured formal expressions, ensuring consistent and verifiable reasoning structures.
3. Analysis Execution: The system must generate and execute precise analysis code that processes validation data according to these formal expressions.
The architecture of GateLens is driven by these fundamental requirements, establishing a systematic pipeline for transforming user queries into analytical results. The system leverages RA to enhance LLMs’ reasoning capabilities and transform user queries into formal relational operations. To support this transformation, GateLens employs domain-specific data schemas that guide its relational modeling and enhance the generation of RA expressions. This approach ensures both accessibility for non-technical stakeholders and the precision required for automotive software validation.
Figure 2: GateLens top-level architecture: The system processes high-level queries from the end user, generates the necessary data manipulation code using the enhanced reasoning layer with the help of RA, executes it, and outputs the result table as a decision-support resource.
3.1 System Overview
The primary objective of GateLens is to generate executable code that performs precise test data analysis based on user queries. The system’s workflow consists of two main phases: query interpretation and code generation. As illustrated in Figure 2, GateLens first processes user queries through an LLM agent that translates natural language inputs into formal RA expressions. This translation incorporates a detailed relational model of the test data, ensuring precise specification of automotive domain concepts. The resulting RA expressions serve as a pivotal intermediate representation that is more transparent to both LLM agents and humans. In the second phase, these formal expressions are passed to the coder agent, which generates executable desirable code, such as SQL or Python code, to perform the required analysis on the test data and produce results.
3.2 Core Components
The system architecture consists of two primary components that work in tandem to transform natural language queries into executable code: the query interpreter agent and the coder agent. The query interpreter agent first translates user queries into RA expressions, providing a structured framework for analytical reasoning. The coder agent then converts these formal expressions into executable code, completing the transformation pipeline. This two-stage approach ensures both analytical precision and efficient implementation, where the prompt engineering flow and prompt structure are presented in Figure 3.
3.2.1 Query Interpreter
The query interpreter agent is responsible for converting user queries into formal RA expressions, providing a precise framework for analytical reasoning. Before initiating this translation, the agent consults the knowledge base, comprising the data schema and domain-specific context. The data schema provides a detailed understanding of the dataset, including its relational modeling, field descriptions, data types, and enriched metadata capturing domain-specific acronyms and terminology mappings. This glossary-enhanced schema is injected into the prompt context, enabling accurate resolution of domain-specific terms into formal schema attributes during RA construction.
To handle imprecise queries without compromising data privacy or exceeding LLM context windows, GateLens employs a selective strategy for exposing categorical information. For low-cardinality attributes (e.g., fields with only a few distinct categories such as test status), all valid options are explicitly enumerated within the schema metadata. In contrast, for high-cardinality attributes, actual database values are never exposed; instead, the schema specifies structural patterns or expected formats (e.g., standard prefixes or identifier conventions). This hybrid design keeps the prompt context compact and privacy-preserving while providing the LLM with sufficient context to generate accurate filtering conditions.
Using this information, the agent verifies whether the query is relevant and within the scope of the dataset [manik2021out]. This validation step ensures that only supported and meaningful queries are processed, improving both accuracy and efficiency. Once the query is confirmed to be in scope, the agent leverages the data schemas to interpret and decompose the query into formal RA expressions.
The agent’s primary function is to map natural language queries into formal RA expressions, enhancing LLM reasoning through structured decomposition [khot2022decomposed]. This approach extends traditional Chain-of-Thoughts (CoT) [wei2022chain] reasoning by constraining the model to think within a formal system framework [zhang2023igniting]. Instead of generating free-form solutions, the agent must express analytics using standard relational algebra notations through a limited set of standard operations: selection, projection, union, set difference, cartesian product, and rename as basic operations, as well as derived operations such as join, intersection, and division and complemented by aggregation functions like average, minimum, maximum, sum, and count.
By limiting operations to this standard set, the agent effectively handles ambiguous queries through formal translation, ensures technical precision, and prevents deviation from analytical requirements. The formal nature of RA enables query optimization, which the agent incorporates by prioritizing data reduction operations early in the expression chain. This optimization strategy involves applying filters first, then performing expensive operations on the reduced dataset, thereby minimizing processing time and resource utilization.
The translation to RA offers two significant advantages. First, it makes the analytics more transparent in technical terms, allowing for clear interpretation and validation of the reasoning process. Second, it ensures that every solution generated is precisely defined and feasible for implementation, preventing the agent from proposing impractical or undefined analytical approaches.
3.2.2 Coder
The coder agent is responsible for generating executable code from given RA expressions. Upon receiving an RA expression, the agent follows precise instructions to produce code that delivers the final analytical results. This capability allows the agent to generate complete, self-contained code at once, eliminating the need for step-by-step generation and execution phases.
To ensure the generated code meets quality standards, we designed prompts that specifically instruct the LLM to include essential validation mechanisms. The prompts explicitly require data type validation instructions for numerical and categorical field handling, null value checking procedures to maintain data integrity, validation of numerical operations, and validation of join conditions to ensure proper matching of key columns between tables.
By generating the entire code in a single pass rather than through iterative refinement, the agent significantly reduces processing overhead, system response time, and resource consumption while minimizing potential errors that could arise from multiple execution steps. This streamlined yet precise approach ensures both efficiency and reliability in the analysis pipeline, fully aligned with the formal rigor of RA expressions and improving overall system responsiveness to user queries.
Figure 3: Overview of the prompt engineering flow and prompt structure within the GateLens system architecture.
3.3 Data Handling
A key architectural decision in GateLens is its indirect interaction with test data. Rather than exposing sensitive test data directly to LLM agents, which could raise privacy concerns [boudewijn2023privacy] and exceed input context limitations, the system operates on data schemas and relational models. This approach serves multiple critical purposes:
  • Privacy Protection: Sensitive automotive test data remains secure within the organization’s infrastructure
  • Scalability: The system can handle large-scale test datasets that would exceed LLM context windows
  • Knowledge Integration: Data schemas and relational models serve as an essential knowledge base, providing necessary structural understanding without raw data exposure
The final execution of the generated code runs on the test data in the target environment, maintaining data privacy while delivering precise analytical results.
4 Experimental Evaluation
In this section, we aim to answer the following research questions:
  1. RQ1:
    How effectively does GateLens address user queries and deliver accurate results across various query categories?
  2. RQ2:
    How robust is GateLens in handling out of scope queries and imprecise queries?
  3. RQ3:
    How does the RA reasoning procedure contribute to the overall performance of GateLens?
  4. RQ4:
    To what extent does RA-based reasoning eliminate the need for in-context learning?
4.1 Experimental Setup
To address the research questions introduced in Section 4, we designed and conducted extensive experiments to evaluate the performance of GateLens.
The experimental data comprises two distinct benchmarks. The first benchmark consists of 50 queries designed with the assistance of release engineers, quality engineers, and verification engineers. To assess GateLens’s performance across a spectrum of query complexities, these queries are categorized into four difficulty levels. The four levels of query difficulty are defined as follows:
Level 1 Simple queries involving a single operation such as filtering or sorting.
Level 2 Queries combining two or three basic operations, such as multiple filtering followed by sorting.
Level 3 Queries involving more than three operations, potentially including grouping and aggregating.
Level 4 Complex queries requiring multiple advanced operations beyond basic filtering and sorting, such as grouping and aggregating for statistical calculations.
The second benchmark is derived from real-world user queries collected from production logs at our partner company. These queries were sourced from the historical logs of an agentic system that employed a well-established tabular data reasoning approach combining CoT [wei2022chain] prompting with Self-Consistency (SC) [wang2022self]. This system was used in production to support software release analytics, and the collected queries reflect a wide range of user roles, query types, and domain-specific requirements. While this system performs effectively in many scenarios, its limitations become apparent as the range of roles and users expands, leading to a significant diversification of queries. This broader query diversity exposes the system’s reliance on few-shot examples, making it less capable of handling highly complex, ambiguous, or ill-defined queries that require greater flexibility and adaptability. Nevertheless, this preliminary system played a critical role in data collection for GateLens by providing query logs used to develop and validate our approach. From these logs, we filtered out near-duplicates and selected 244 frequently repeated unique queries, which we then organized into eight functional categories based on their purposes. The ground truth for the 244 real-world queries was established through a two-stage manual annotation process: two domain experts independently solved an initial subset of 20 queries to align on interpretation and expected outputs, after which the remaining queries were divided and annotated following the agreed-upon criteria, with periodic cross-checks to ensure consistency.
In order to assess GateLens’s performance, we run experiments with two large language models (LLMs): GPT-4o, a leading commercial model, and Llama 3.1 70b, a recently released open-source model. We also benchmark GateLens against the CoT+SC [khoee2024gonogo] agentic system currently used to support the company’s release decisions. Our comparative analysis is designed to quantify the improvements introduced by GateLens’s novel architecture.
To address the challenge of handling out-of-scope user queries during real-time interactions, GateLens incorporates an in-scope filtering mechanism as explained in Section 3.2. This mechanism ensures that the system only attempts to process queries that fall within its scope, thereby improving reliability and reducing errors. Performance evaluation focused on two key aspects:
  1. 1.
    Quality of responses: Measured using precision, recall, and F1 Score, which reflect the system’s ability to address relevant queries correctly.
  2. 2.
    Coverage of relevant queries: Ensuring the system does not reject a significant proportion of valid queries, thus maintaining broad applicability.
Additionally, an ablation study is conducted to examine the contribution of the RA reasoning mechanism of GateLens.
In our experiments, the evaluation of system performance is based on the following definitions: A True Positive (TP) occurs when the system produces a result that matches the manually generated ground-truth result, specifically when the final output of the executed code matches the expected ground-truth output. A False Positive (FP) occurs when the system provides an incorrect result, meaning the executed code produces output that differs from the ground-truth. A False Negative (FN) occurs when the system fails to provide any result for a query. True Negatives (TNs) are not applicable since we focus on valid queries producing meaningful output.
Based on these definitions, we calculate Precision, Recall, and F1 scores to assess the performance of the system. Precision ensures that incorrect results are minimized, recall ensures relevant queries are addressed, and the F1 score balances the two to provide an overall assessment of system performance. By relying on a closed-set benchmark with established ground truths, these metrics enable us to rigorously isolate and measure the impact of our architectural choices during lab validation, setting the stage for the real-world industrial evaluation detailed in Section LABEL:sec:lessons.
4.2 Performance in Addressing User Queries (RQ1)
Table 1: Comparison of average token consumption and reduction between CoT+SC and GateLens across four difficulty levels. The evaluation was conducted on the first benchmark, which consists of 50 designed queries with annotated difficulty levels, with both agents utilizing GPT-4o.
Level # Queries GateLens with GPT-4o GateLens with Llama 3.1 70B CoT+SC with GPT-4o CoT+SC with Llama 3.1 70B
Precision Recall F1 Score Precision Recall F1 Score Precision Recall F1 Score Precision Recall F1 Score
1 16 100% 100% 100% 100% 43.75% 60.87% 93.33% 87.5% 90.32% 100% 93.75% 96.77%
2 16 100% 100% 100% 100% 62.5% 76.92% 100% 81.25% 89.66% 92.31% 75% 82.76%
3 12 100% 100% 100% 100% 50% 66.67% 91.67% 91.67% 91.67% 90.91% 83.33% 86.96%
4 6 100% 100% 100% 100% 33% 49.62% 66.67% 66.67% 66.67% 60% 50% 54.55%
Total 50 100% 100% 100% 100% 47.31% 63.52% 87.91% 81.77% 84.57% 85.81% 75.52% 80.26%
Table 2: Performance comparison of GateLens and the CoT+SC system across different categories on the second benchmark, which consists of 244 real-world queries.
Category # Queries GateLens with GPT-4o GateLens with Llama 3.1 70B CoT+SC with GPT-4o
Precision Recall F1 Score Precision Recall F1 Score Precision Recall F1 Score
Column Operations 17 64.7% 64.7% 64.7% 50% 11.76% 19.04% 76.47% 76.47% 76.47%
Complex Multi-Condition Queries 77 86.3% 81.82% 84% 100% 24.68% 39.59% 84.75% 64.94% 73.53%
Conditional Calculations 8 100% 87.5% 93.3% 100% 37.5% 54.55% 87.5% 87.5% 87.5%
Data Filtering 32 89.66% 81.25% 85.25% 90.91% 31.25% 46.51% 86.21% 78.13% 81.97%
Duplicate Removal 78 87.67% 82.05% 84.77% 100% 23.08% 37.5% 75.93% 52.56% 62.12%
Grouping and Aggregation 10 80.0% 80.0% 80.0% 100% 30% 46.15% 83.33% 50% 62.5%
Metadata Queries 13 91.67% 84.61% 88.0% 100% 15.38% 26.66% 46.15% 46.15% 46.15%
Table Generation 9 88.89% 88.89% 88.89% 100% 44.44% 61.53% 88.89% 88.89% 88.89%
Total 244 86.02% 81.14% 83.51% 92.61% 27.26% 41.44% 83.15% 63.52% 70.61%
We conducted experiments to compare the performance of GateLens across the two introduced benchmarks. The first benchmark, consisting of 50 queries categorized by difficulty levels, was used to evaluate and compare the performance of GateLens and CoT+SC. Both systems were tested using GPT-4o and Llama 3.1 70B as their underlying LLMs. The results are summarized in Table 1.
The results demonstrate that GateLens with GPT-4o significantly outperforms GateLens with Llama 3.1 70B, indicating GPT-4o’s superior capability for interpreting and generating RA. Similarly, CoT+SC with GPT-4o outperforms its Llama 3.1 70B variant, with the performance gap growing as query complexity increases. CoT+SC performance declines with query complexity. This underscores the importance of the RA reasoning mechanism in GateLens, which enables effective handling of complex, unstructured queries by decomposing them into logical, structured expressions. Most notably, GateLens with GPT-4o achieved optimal performance on this benchmark, maintaining 100% accuracy across all difficulty levels. This stems from integrating RA reasoning into our framework. By translating queries into RA expressions, GateLens explicitly captures the logical structure of operations, enhancing both the clarity and precision of the generated code. The intermediate RA conversion allows the system to focus on the relevant table operations while filtering out irrelevant elements in the query, greatly enhancing the problem-solving capabilities of the LLM agent.
For the second benchmark, results in Table 2 show that GateLens (GPT-4o) and CoT+SC (GPT-4o) significantly outperformed GateLens with Llama 3.1 70B. This performance disparity is primarily due to the strict code generation requirements of the task, including table filtering, merging strategies, and key-value mapping operations, where GPT-4o demonstrated markedly superior capabilities.
GateLens with GPT-4o outperformed CoT+SC (GPT-4o) across most categories, particularly evident in Metadata Queries (those seeking basic table information). For example, when processing the query ”Give me the list of release candidates,” the CoT+SC system often fails to identify the correct field. A common failure mode in CoT+SC occurred when user queries included typographical errors or incorrect casing in field names, with the system directly using the erroneous fields without correction. GateLens addresses this limitation through its query-to-RA transformation process, which incorporates the database’s relational model, adjusts query fields to match table formats, and can handle fuzzy matching to detect and correct field names, enabling the system to resolve typographical errors and ambiguous queries effectively. This approach improves accuracy and resilience, particularly in real-world scenarios where user queries may not always adhere to strict formatting standards.
Table 3: Comparison of average token consumption and reduction between CoT+SC and GateLens across four difficulty levels. The evaluation was conducted on the first benchmark, which consists of 50 designed queries with annotated difficulty levels, with both agents utilizing GPT-4o.
Level Agent Avg Input Tokens Avg Output Tokens Avg Total Tokens Token Reduction
1 CoT+SC 11,905 186 12,091
GateLens 2,239 420 2,658 $\downarrow$ 78%
2 CoT+SC 13,747 368 14,116
GateLens 2,428 701 3,129 $\downarrow$ 78%
3 CoT+SC 14,726 441 15,168
GateLens 2,432 698 3,130 $\downarrow$ 79%
4 CoT+SC 17,103 452 17,555
GateLens 2,505 847 3,352 $\downarrow$ 81%
In addition, compared to the CoT+SC solution, GateLens uses significantly fewer tokens due to its effective use of RA as the intermediate representation and its zero-shot architecture. In Table 3, we compare the token usage of both systems across different difficulty levels. Notably, while the CoT+SC approach requires increasingly massive input contexts, primarily to accommodate few-shot examples, GateLens maintains a highly compact input footprint. As the query difficulty increases, the token reduction becomes even more significant, demonstrating the efficiency and scalability of the RA-based approach in handling complex queries.
RQ1 findings: GateLens with GPT-4o achieved 100% F1 score on the first benchmark across all difficulty levels and 83.51% F1 score on the second benchmark with 244 real-world queries. It outperformed CoT+SC (GPT-4o) by approximately 13 percentage points on real-world queries, indicating that the RA reasoning mechanism effectively addresses both complex logical operations and real-world noise, such as typos and ambiguous field names.
4.3 Robustness: Handling Out of Scope and Imprecise Queries (RQ2)
To assess the robustness of our approach in handling diverse user queries under real-world conditions, we conducted further experiments focusing on filtering out-of-scope queries as well as processing imprecise queries. For this purpose, the data analysis team at our industrial partner company manually selected 37 out-of-scope queries and 50 imprecise queries from the historical logs of the first-generation system, which are used to perform targeted evaluations.
Out-of-scope queries are those that cannot be meaningfully answered using the available data. For example, a query like ”What is the most beautiful truck?” requires subjective judgment and cannot be resolved through database operations; it should be identified and filtered as out of scope. On the other hand, imprecise queries are those that can be answered using the database but contain ambiguous or inexact terms. For instance, a query such as ”Find some trucks for cases that are NOK” is considered imprecise because while it seeks truck names where test results are ”NOK” (failed), it uses ambiguous terminology - referring to ”trucks” instead of the actual database field ”name”, and mentions ”NOK” without specifying the ”test_result” field. Such imprecise queries require mapping informal language to precise database fields and conditions for proper execution.
4.3.1 Handling Out of Scope Queries
We compared GateLens with other models; the results can be found in Table 4. The results demonstrate that GateLens with GPT-4o achieved the best performance, particularly in terms of precision, which is approximately 40% higher than other models, indicating GateLens’ ability to avoid generating incorrect results.
The superior precision of GateLens with GPT-4o can be attributed to two key aspects of its design. First, its robust filtering mechanism ensures that out-of-scope queries are identified and excluded early in the processing pipeline, preventing irrelevant results. Second, the conversion of raw natural language queries into structured RA expressions enables the model to isolate and capture task-relevant components of a query. This structured approach considerably decreases erroneous outcomes and enhances the model’s ability to handle complex and diverse query formulations in real-world scenarios.
Table 4: Model comparison for out-of-scope queries.
Model Precision Recall F1 Score
GateLens with GPT-4o 92.5% 100% 96.10%
GateLens with Llama 3.1 70B 52.94% 97.30% 68.57%
CoT+SC with GPT-4o 51.10% 89.19% 64.97%
CoT+SC showed significantly lower precision due to the variability of real-world queries and the inconsistency of user narratives, which often contain a mix of relevant and irrelevant content. This variability increases uncertainty and poses challenges for models that struggle to identify task-relevant information. Although all models demonstrated high recall, this did not translate into accurate processing.
4.3.2 Handling Imprecise Queries
To further assess the robustness of our approach, we evaluated its performance in handling imprecise queries, which posed two primary challenges. First, some queries are informal and conversational in style, appearing unrelated to data analysis but actually carrying relevant intent. Second, many queries referred to fields using terms differing from the column headers.
Table 5: Model comparison for imprecise queries.
Model Precision Recall F1 Score
GateLens with GPT-4o 92.86% 78% 84.78%
GateLens with Llama 3.1 70B 92.86% 26% 40.63%
CoT+SC with GPT-4o 90% 36% 51.43%
The results of these experiments are presented in Table 5. As shown, GateLens with GPT-4o demonstrates the best overall performance. In terms of precision, all methods performed relatively well, indicating that when results are generated, they are likely to be correct. However, our method significantly outperformed the others in recall, highlighting its ability to handle a larger portion of the imprecise queries. As a result, GateLens with GPT-4 achieved a substantially higher F1 score compared to other methods, demonstrating that it not only processes most queries but also produces accurate results for them.
The observed performance gap between GateLens with GPT-4o and the other models can be attributed to their inherent limitations. Specifically, the Llama 3.1 70B model struggled to interpret user queries that deviated from the exact column header descriptions in the database schema. In such cases, Llama 3.1 70B often converted only the clearly defined parts of the query into RA, leading to incomplete execution and reduced accuracy. On the other hand, CoT+SC exhibits low recall, as it is highly susceptible to confusion by ambiguous query elements. This causes CoT+SC to frequently generate incorrect code that fails execution, significantly lowering its recall rate.
RQ2 findings: GateLens demonstrates high robustness in real-world conditions. It effectively filters out-of-scope inputs, achieving approximately 40% higher precision than the baseline system. Furthermore, it handles imprecise or informal queries with superior performance, more than doubling the recall (78% vs. 36%). This confirms that the system can manage ambiguous user inputs without sacrificing the accuracy of the generated code.
4.4 Effectiveness of the RA module (RQ3)
(a) The first benchmark with annotated difficulty levels.
(b) The second benchmark with real-world user queries
Figure 4: Comparison of the original method and the method without the RA module across different datasets.
To evaluate the impact of the RA module that converts user queries into RA expressions, we conducted experiments by removing the RA module from the framework and comparing the results to the original system. The outcomes, shown in Figure 4, demonstrate significant performance degradation across both benchmarks when operating without the RA module.
In the first benchmark, performance declined most notably for Level 4 queries, showing a drop exceeding 27%. These queries, which involve advanced operations like grouping, aggregating, and statistical calculations, demonstrated that RA translation is particularly crucial for handling queries with multiple, intricate operations. Similarly, the second benchmark shows decreased performance in complex tasks such as multi-condition filtering, duplicate data removal, and table generation, further emphasizing RA’s effectiveness in managing complex database operations.
The RA module maintained consistent performance for simpler queries, demonstrating its versatility across varying complexity levels. By transforming natural language into precise, logical representations, the RA module serves as a key bridge between user intent and code execution. This translation process enables the code generator to produce accurate, efficient executable code for data analysis tasks.
RQ3 findings: The RA module is a critical component for handling query complexity. Removing it results in substantial performance degradation (over 27% for complex queries), confirming that translating natural language to RA provides necessary structural guidance for accurate code generation.
4.5 Role of Few-Shot Examples (RQ4)
To investigate the effect of including few-shot examples in prompts, we conducted experiments by varying the number of examples provided to both GateLens and CoT+SC. This experiment is performed on the first benchmark containing 50 designed queries, with results illustrated in Figure 5.
The results demonstrate that GateLens relies heavily on its RA translation process, achieving 100% F1 even in a 0-shot setting without any examples. Interestingly, when a small number of examples are added, GateLens becomes slightly biased toward them, leading to a minor degradation in performance (dropping to 95.92% at 2 examples). However, it regains optimal performance at 3 examples and remains at 100% thereafter. In contrast, CoT+SC’s performance heavily depends on in-context examples, achieving only 42% F1 with 2 examples and showing suboptimal results without sufficient few-shot examples. Performance improved steadily with more examples, reaching 84.57% at 50 examples.
Figure 5: Comparison of GateLens against CoT+SC across different numbers of few-shot examples.
RQ4 findings: GateLens does not require many-shot examples to achieve high performance, delivering optimal results in a 0-shot setting. This contrasts with CoT+SC, which depends on carefully curated examples. This independence from few-shot examples makes GateLens more efficient for real-time applications by reducing context size and computational overhead.
5 Qualitative Assessment Examples
Table 6: An example of an input query and its corresponding relational algebra expressions translated using a domain-specific database schema, along with the executable code generated from the relational algebra and the resulting output produced when executed over the database.
Stage Content
Input Query Compute the relative frequency of not run tests for each RC and function
Relational Step 1: $R0\leftarrow\sigma_{\text{result\_status}=\text{"NotRun"}}(\text{df})$
Algebra Step 2: $R1\leftarrow\gamma_{\text{releasecandidate},\text{function},\text{countNotRun}\rightarrow\text{count}}(R0)$
Step 3: $R2\leftarrow\gamma_{\text{releasecandidate},\text{function},\text{totalcount}\rightarrow\text{count}}(\text{df})$
Step 4: $R3\leftarrow R1\bowtie_{\text{R1.releasecandidate}=\text{R2.releasecandidate}\wedge\text{R1.function}=\text{R2.function}}R2$
Step 5: $\text{Result}\leftarrow\gamma_{\text{R3.releasecandidate},\text{R3.function},\text{relativeFrequency}\rightarrow\frac{\text{R3.count}}{\text{R3.totalcount}}}(R3)$
Python Code